OpenStudy (anonymous):

perimeter ? L-14m w-4m

hartnn (hartnn):

is it a rectangle ?

hartnn (hartnn):

if its a rectangle , then perimeter = 2(L+w) = 2(14+4) = 2(28) =56 ok?

OpenStudy (anonymous):

14+14+4+4=36

hartnn (hartnn):

oops, sorry 2(18) = 36